# Weipu Chen, Zhuangzhuang He, Fei Liu*

### v1

@@ I Introduction
Therefore, we propose an Adaptive Ensemble Learning (AEL) for denoising recommendation. Inspired by SparseMoE (Sparse Mixture-of-Experts) [26], AEL consists of three experts and a sparse gating network. Specifically, we construct the experts of the denoising module in a novel approach and introduce an adaptive ensemble module to fundamentally address adaptability insufficiency. For C1, we propose a novel method that stacks components to construct sub-recommenders. It not only reduces the model complexity but also ensures sub-recommender diversity. For C2, we creatively introduce a sparse gating network, which can analyze the performance of each sub-recommender and generate suitable dynamic weight distributions for current input. In summary, our contributions are as follows:
@@ III-A Experimental Settings
Implementation Details. The hidden dimensions of sub-AEs are set to 12, 48, and 128, respectively. The number of experts $k$ activated per input is set to 2. The coefficients of $L_{importance}$ and $L_{Load}$, $w_{i}$ and $w_{l}$, are set to 1e-2. For other baselines, we follow the hyperparameter settings suggested in  [9, 10].
@@ III-C Extensive Effectiveness Analysis
Hyperparametric Sensitivity Analysis. The number of experts $k$ selected is the most critical hyperparameter, as it controls how to synthesize the denoising capacity for current input. We experiment with all the feasible cases of $k$ and show the results in the Table V. We observed that AEL performs well across three datasets in all cases of $k$. Although AEL is slightly worse when $k=2$ than when $k=3$ in Recall@5 on the Adressa dataset, it shows overall surpass in other cases. That reason is that for each interaction vector $x_{u}$, activating two suitable parent-AEs out of three can provide a sufficient diversity of denoising capacities compared to activating one, and also eliminate the effect of unnecessary experts. Setting $k=2$ helps prevent sending inappropriate samples to a parent-AE that inadequately or overly denoise them.
Aggregation Method Comparison. We conduct experiments to explore how the aggregation methods affect AEL performance. We compare other ensemble learning methods, including Bayesian Model Averaging and Averaging, to our spare gating network. According to the results shown in Fig. 4, our aggregation method outperforms others across all datasets, especially on the MovieLens dataset.

### v3

# Weipu Chen, Zhuangzhuang He, Fei Liu School of Computer Science and Information Engineering

@@ I Introduction
Therefore, we propose an Adaptive Ensemble Learning (AEL) for denoising recommendation. Inspired by SparseMoE (Sparse Mixture-of-Experts) [26], AEL consists of three experts and a sparse gating network. Specifically, we construct the experts of the denoising module in a novel approach and introduce an adaptive ensemble module to fundamentally address adaptability insufficiency. For C1, we propose a novel method that stacks components to construct sub-recommenders. It not only reduces the model complexity but also ensures sub-recommender diversity. For C2, we creatively introduce a sparse gating network, which can analyze the performance of each sub-recommender and generate suitable dynamic weight distributions for current input.
In summary, our contributions are as follows:
@@ III-A Experimental Settings
Implementation Details. The hidden dimensions of sub-AEs are set to 12, 48, and 128, respectively. The number of experts $k$ activated per input is set to 2. The coefficients of $L_{importance}$ and $L_{Load}$, $w_{i}$ and $w_{l}$, are set to 1e-2. For baselines, we follow the hyperparameter settings suggested in  [9, 10].
@@ III-C Extensive Effectiveness Analysis
Hyperparametric Sensitivity Analysis. The number of experts $k$ selected is the most critical hyperparameter, as it controls how to synthesize the denoising capacity for current input. We test all feasible $k$ values, as shown in Table V. We observed that AEL performs well across three datasets in all cases of $k$, with particularly strong performance when $k=2$. That reason is that for each interaction vector $x_{u}$, activating two suitable parent-AEs offers sufficient denoising diversity while avoiding unnecessary experts. Setting $k=2$ helps prevent sending inappropriate samples to a parent-AE that inadequately or overly denoise them.
Aggregation Method Comparison. We conduct experiments to explore how the aggregation methods affect AEL performance. We compare other ensemble learning methods, including Bayesian Model Averaging (BMA) and Averaging, to our spare gating network. According to the results shown in Fig. 4, our aggregation method outperforms others across all datasets, especially on the MovieLens dataset.